OpenStudy (destiny.toddd):

If f(x) = x2 and g(x) = 1 2x + 3 , find g(f(-1))

rebeccaxhawaii (rebeccaxhawaii):

Plug in -1 into the f(x) equation

Mehek (mehek14):

First you have to find (f-1) = \(-1^2\)

Mehek (mehek14):

After that, the value of (f-1), you put for x in g(x)

OpenStudy (destiny.toddd):

1/2x+3

Mehek (mehek14):

that's for g right?

OpenStudy (destiny.toddd):

yes , 1 was over 2x+3

Mehek (mehek14):

\(\dfrac{1}{2}x+3\) like that?

OpenStudy (destiny.toddd):

no 1 over 2x+3

Mehek (mehek14):

\(\dfrac{1}{2x+3}\)

Mehek (mehek14):

is that how it looks?

OpenStudy (destiny.toddd):

yes

Mehek (mehek14):

Alright first find f(-1) = \(-1^2\) = ?

OpenStudy (phi):

do you get the idea that f(x) is the name of a "rule" and in this case the rule is x^2 (which means x*x) and that means: if you are old x is -1, then using the rule f(x) means do -1*-1 and that is what f(-1) is

OpenStudy (phi):

once you know what f(-1) is (as a number) you can do g( whatever number f(-1) is) and g is a different rule.